# On a class of interdiction problems with partition matroids: complexity and polynomial-time algorithms

### v3

@@ 4 Conclusion
In this study, we explore the theoretical computational complexity of a partition matroid interdiction (PMI) problem. The problem can be viewed as a zero-sum game between a leader and a follower, whose feasible sets are induced by partition matroids over the same weighted ground set. The leader blocks an independent set of its partition matroid so as to minimize the weight of the maximum weight independent set in the follower’s matroid. In particular, the follower is not allowed to use the elements blocked by the leader. From a practical perspective, the PMI problem can also be interpreted as a zero-sum game between multiple leaders and multiple independent followers. In particular, the leaders and the followers operate over two different partitions of the ground set and each decision-maker is subject to a single budget constraint with unit costs.
@@
As an additional interesting finding, we demonstrate that the PMI problem corresponds to minimizing a monotone submodular function subject to cardinality or matroid constraints. In particular, it turns out that, under a mild assumption, the problem with a single capacity constraint for the follower can be solved exactly with a simple greedy algorithm for the leader. However, if the follower has more than one capacity constraint, then the greedy algorithm fails to provide any constant-factor approximation for the PMI problem.
In conclusion, future research could focus on the following ideas. On the one hand, it is interesting to explore whether there exist polynomial-time constant-factor approximation algorithms for the considered PMI problem, given that the numbers of capacity constraints are not fixed. We believe that these algorithms could leverage the submodularity and monotonicity of the leader’s objective function. However, while the problem of minimizing a submodular function without constraints can be solved in polynomial time [28], there are only few studies that consider submodular minimization with multiple constraints; see, e.g., [29, 32]. Alternatively, potential approximation algorithms could draw on the ideas of Chen et al. [11], who designed polynomial-time approximation algorithms for a more general multidimensional knapsack interdiction problem but with fixed numbers of budget constraints.
@@
$\psi(\mathbf{x})=\sum_{k\in K_{f}}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}),$
@@
$\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}):=\max_{\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}}\Big{\{}\boldsymbol{\beta}^{\mbox{\tiny(k}\mbox{\tiny)}\top}\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}:\;\mathbf{0}\leq\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}\leq\mathbf{1}-\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}},\;\mathbf{1}^{\top}\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}\leq f_{k}\Big{\}},$ (25)
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Clearly, the increase of $\mathbf{x}$ can only reduce the follower’s feasible sets in (25) and, therefore, it cannot increase the leader’s objective function value, $\psi(\mathbf{x})$. Hence, $\psi(\mathbf{x})$ is monotone.
@@
$\psi(\mathbf{x}^{\prime}+\mathbf{e}_{i})-\psi(\mathbf{x}^{\prime})=\psi_{k}(\mathbf{x}^{\prime\mbox{\tiny(k}\mbox{\tiny)}}+\mathbf{e}_{i})-\psi_{k}(\mathbf{x}^{\prime\mbox{\tiny(k}\mbox{\tiny)}})=\begin{cases}\beta_{j_{k}(\mathbf{x}^{\prime},f_{k}+1)}-\beta_{i},\mbox{ if }i\geq j_{k}(\mathbf{x}^{\prime},f_{k}),\\ 0,\mbox{ otherwise.}\end{cases}$

### v4

@@ 4 Conclusion
In this study, we explore the theoretical computational complexity of a partition matroid interdiction (PMI) problem. The problem can be viewed as a zero-sum game between a leader and a follower, whose feasible sets are induced by partition matroids over the same weighted ground set. The leader blocks an independent set of its partition matroid so as to minimize the weight of the maximum weight independent set in the follower’s matroid. In particular, the follower is not allowed to use the elements blocked by the leader. From a practical perspective, the PMI problem can also be interpreted as a zero-sum game between multiple leaders and multiple independent followers. The leaders and the followers operate over two different partitions of the ground set and each decision-maker is subject to a single budget constraint with unit costs.
@@
As an additional interesting finding, we demonstrate that the PMI problem corresponds to minimizing a monotone non-increasing submodular function subject to matroid constraints. In particular, it turns out that, under a mild assumption, the problem with a single capacity constraint for the follower can be solved exactly with a simple greedy algorithm for the leader. However, if the follower has more than one capacity constraint, then the greedy algorithm fails to provide any constant-factor approximation for the PMI problem.
In conclusion, future research could focus on the following ideas. On the one hand, it is interesting to explore whether there exist polynomial-time constant-factor approximation algorithms for the considered PMI problem, given that the numbers of capacity constraints are not fixed. We believe that these algorithms could leverage the submodularity and monotonicity of the leader’s objective function. However, while the problem of minimizing a submodular function without constraints can be solved in polynomial time [28, 40], there are only few studies that consider submodular minimization with multiple constraints; see, e.g., [29, 32]. Alternatively, potential approximation algorithms could draw on the ideas of Chen et al. [11], who designed polynomial-time approximation algorithms for a more general multidimensional knapsack interdiction problem but with fixed numbers of budget constraints.
@@
$\psi(\mathbf{x})=\sum_{k\in K_{f}}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}}),$
@@
$\psi_{k}(\mathbf{x}^{\text{\tiny({k})}}):=\max_{\mathbf{y}^{\text{\tiny({k})}}}\Big{\{}\boldsymbol{\beta}^{\text{\tiny({k})}\top}\mathbf{y}^{\text{\tiny({k})}}:\;\mathbf{0}\leq\mathbf{y}^{\text{\tiny({k})}}\leq\mathbf{1}-\mathbf{x}^{\text{\tiny({k})}},\;\mathbf{1}^{\top}\mathbf{y}^{\text{\tiny({k})}}\leq f_{k}\Big{\}},$ (26)
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Clearly, the increase of $\mathbf{x}$ can only reduce the follower’s feasible sets in (26) and, therefore, it cannot increase the leader’s objective function value, $\psi(\mathbf{x})$. Hence, $\psi(\mathbf{x})$ is monotone non-increasing.
@@
$\psi(\mathbf{x}^{\prime}+\mathbf{e}_{i})-\psi(\mathbf{x}^{\prime})=\psi_{k}(\mathbf{x}^{\prime\text{\tiny({k})}}+\mathbf{e}_{i})-\psi_{k}(\mathbf{x}^{\prime\text{\tiny({k})}})=\begin{cases}\beta_{j_{k}(\mathbf{x}^{\prime},f_{k}+1)}-\beta_{i},\mbox{ if }i\geq j_{k}(\mathbf{x}^{\prime},f_{k}),\\ 0,\mbox{ otherwise.}\end{cases}$